# Memento 2: Learning by Stateful Reflective Memory

### v2

@@ Abstract
We present a theoretical study for continual and experiential learning in Large Language Model (LLM) agents integrating episodic memory with reinforcement learning. We recognise that reflection, the ability of an LLM to examine past outcomes and adjust future actions, serves as the key cognitive mechanism for continual adaptation without LLM fine-tuning. This challenges the traditional separation between training and testing in machine learning. Our empirical studies on Memento [zhou2025memento] and case-based reasoning LLMs [guo2025optimizing, guo2024ds] indicate that episodic-experience-driven reflection enables continual adaptation in open-ended, long-horizon tasks, motivating the need for a formal characterisation of the underlying learning and control mechanisms. In this paper, we fill this gap by introducing the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This frames reflective memory as a control theoretic object amenable to classical reinforcement learning analysis. We instantiate this Read–Write Reflective Learning by integrating retrieval into soft policy iteration, with guaranteed convergence. We formally show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning, retrieval-augmented generation with principled reinforcement learning, and provides a rigorous mathematical basis for building reflective, memory-embeded LLM agents capable of continual general-purpose learning.
Figure 1: Learning by Stateful Reflection with Memory realises continual learning without fine-tuning LLMs by iteratively reading from (policy improvement) and writing to (policy evaluation) an evolving episodic memory that drives the agent’s online adaptation. Empirically, this Read-Write Learning paradigm has shown strong effectiveness across diverse domains, including software testing [guo2025optimizing], automated data science [guo2024ds], and deep research agents [zhou2025memento], demonstrating that memory-driven reflection can endow LLM agents with genuine self-improving capabilities.
@@ 1 Introduction
Modern machine learning is about learning from experience [turing1996intelligent, silver2025welcome]; both supervised learning and reinforcement learning can be viewed as processes of optimising a parameterised model through experience, either from human labelled examples or data-driven feedback. The objective is to learn a mapping from inputs to outputs, or in the case of reinforcement learning, from states to actions, such that the resulting function generalises effectively to new, unseen situations [shalev2014understanding, valiant1984theory]. Formally, the learner assumes that the target function is constrained by a family of parameterised models, for instance, neural networks, where the parameters determine the precise form of the mapping.
@@
In supervised learning, these parameters are typically adjusted by minimising a loss function defined over labelled examples, whereas in reinforcement learning (RL), the supervision signal is indirect: instead of labelled outputs, the agent interacts with an environment, receives rewards, and adjusts its parameters to maximise the expected cumulative return [SuttonBarto2018]. In both cases, the parameters are updated by backpropagation, a gradient-based optimisation algorithm that propagates errors through the network to refine the model [rumelhart1986learning].
@@
Despite their substantial empirical successes, current machine learning paradigms suffer from fundamental inefficiencies in sample complexity. Learning via backpropagation and gradient-based optimisation typically requires an enormous number of training examples, whether labelled data in supervised learning or interaction trajectories in reinforcement learning, before competent behaviour emerges [goodfellow2016deep, guo2025deepseek]. From a theoretical perspective, this reflects the difficulty of identifying generalisable structure in high-dimensional, non-linear hypothesis classes when learning is driven almost entirely by statistical correlations [shalev2014understanding]. In the absence of explicit access to semantic abstractions, compositional structure, or task-level meaning, learning algorithms must recover regularities indirectly through averaging over large numbers of samples, leading to intrinsically high sample complexity.
@@
The training of large language models continuously exemplify this frustration. Although massive pre-training endows them with broad linguistic and reasoning abilities, adapting them to specific tasks or domains still has to resort to additional fine-tuning, such as supervised fine-tuning (SFT) or reinforcement learning with human feedback (RLHF) [ouyang2022training], each demanding further large quantities of curated data or human annotations [lightman2023let].
@@
In contrast, humans leverage semantic representations, episodic memory, and reflective reuse of experience to adapt with far fewer examples [GershmanDaw2017EpisodicRL]. This contrast suggests that learning mechanisms beyond purely statistical optimisation, based on stateful memory and semantic reflection, may offer a principled path to lower complexity learning without large scale gradient updates.
@@
    In-context learning. As early as GPT-3, researchers observed that LLMs could learn from a few human-provided examples directly within a prompt, a process known as in-context learning, where a few labelled examples guide behaviour on new tasks [brown2020language] (Figure 2 (a)). For instance, given a few examples of English-to-French translations in the prompt, GPT-3 can correctly translate a new English sentence into French without any parameter updates.
@@
    Feedback-driven reflection. Subsequent studies showed that given a specific unknown task, LLMs could refine their outputs through interaction with the environment, adjusting responses based on external feedback [Yao2022ReAct] (Figure 2 (b)). For instance, in code generation, the LLM proposes a function, runs tests to observe errors or runtime feedback, and iteratively refines the code to reach a correct solution.
@@
    Internal reasoning. More recently, LLM models have displayed the ability to perform internal reasoning via chain-of-thought processes before generating final outputs [Wei2022CoT] (Figure 2 (c)). For instance, when solving a math word problem, the model explicitly writes intermediate reasoning steps (e.g., decomposing a problem into smaller parts) before computing the final numerical answer.
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection such as information gathering or internal computation to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [brown2020language, guo2025deepseek], but also in multimodal LLMs [driess2023palm] and non-language transformers [hollmann2025accurate, yu2025self].
@@
In our earlier empirical work [guo2024ds, guo2025optimizing, zhou2025memento], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [guo2024ds], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [zhou2025memento] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [guo2025optimizing], automated data science [guo2024ds, Grosnit2024KolbBasedEL], and deep research [zhou2025memento], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
In this paper, we further provide a theoretical analysis and formation and leverage the capability of generalised reflection to develop a new paradigm for sample-efficient learning from experience without relying on data-intensive model training. Instead of updating model parameters, the proposed learning process operates through intermediate reflection, driven by relevant past experiences retrieved from an episodic memory [GershmanDaw2017EpisodicRL], as shown in Fig 1. This framework can unify diverse sources of experience, whether originating from internal reasoning (e.g., chain-of-thought), interactions with the environment (e.g., ReAct [Yao2022ReAct]), or human-provided labelled data, under a single, memory-based learning mechanism.
@@
Within this formulation, learning naturally emerges from the process of retrieval: policy improvement is achieved through more effective retrieval, while policy evaluation is realised through writing experiences back to memory. This mechanism closely mirrors the principles of episodic control and reinforcement learning in humans [GershmanDaw2017EpisodicRL, fountas2024human]. Consequently, such learning can be efficiently implemented using retrieval-augmented generation architectures equipped with large-scale memory systems [Borgeaud2022RETRO, Lewis2020RAG].
A key desire property is its self-learnability. With a frozen LLM [^1] While LLM parameters may evolve over time through offline updates given adequate data, such updates typically occur at a much slower timescale. Our approach does not rule out parameter learning, but instead addresses the intermediate regime in which adaptation must occur without immediate model updates. as its reasoning and reflection core, the agent improves through trial and error when data is limited and later reuses and combines past knowledge as experience grows. It balances exploitation (using known cases), exploration (targeted search in uncertain areas), and discovery (seeking new ones) through adaptive memory operations, achieving continual learning without parameter updates.
@@ 2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [yao2023tree] and Graph of Thoughts [besta2024graph], have demonstrated the potential of structured reflection to enhance LLM reasoning. From a different perspective, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [guo2024ds, guo2025optimizing, zhou2025memento]. To formalise this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which models continual learning as a sequence of reflective interactions with memory.
@@
Figure 1 illustrates the overall workflow of an SRDP. We begin by formally defining episodic memory as a structured repository of past interaction episodes, consistent with the classical notion of episodic memory in cognitive psychology [tulving1983elements].
@@
SRDP generalises and unifies several existing learning paradigms. Compared with case based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [guo2024ds, AamodtPlaza1994], SRDP embeds retrieval within a sequential decision making framework. This explicitly models temporal dependencies, memory evolution, and long horizon effects under a Markovian assumption, rather than treating cases as isolated problem instances.
In contrast to retrieval augmented generation (RAG), which conditions generation on static external documents or databases [Borgeaud2022RETRO], SRDP treats memory as a dynamic and stateful component. Episodic memory is continually updated through interaction with the environment and grounded in evaluative feedback, allowing retrieval to reflect the agent’s own experience rather than fixed external knowledge. From the perspective of memory based reinforcement learning, SRDP replaces parameter level updates with reflective memory updates mediated by an LLM. Policy improvement is achieved through structured recall and reflection over episodic memory rather than gradient based optimisation of model parameters [ramani2019short]. This shift enables continual adaptation while keeping the underlying LLM fixed, positioning reflection as the primary learning mechanism.
This formulation is also strongly inspired by neuroscientific findings on the role of the hippocampus in episodic memory and reflective cognition [hami2025, McClelland1995CLS]. In humans and animals, episodic recall supports flexible reasoning and planning by replaying past experiences, an ability mirrored by the reflective cycles of SRDP. Thus, the Stateful Reflective Decision Process provides both a computational and cognitive bridge: it formalises how memory and reflection together enable continual, self-improving intelligence without explicit retraining.
Experimentally, it has been demonstrated that this approach enables learning without fine-tuning across various agentic tasks, including data-science agents [guo2025optimizing], software-testing agents [guo2024ds], and deep-research agents [zhou2025memento].
@@ 3 A Model for Reflective Process over Episodic Experience
Mathematically, the core idea is to extend the Markov Decision Process [Puterman1994] with an intermediate reflection stage, where the agent retrieves relevant experiences from external memory before generating the final output, enabling continual and memory-based learning.
@@ 3.1 Preliminary: Markov Decision Process
To make self-contained, we begin by reviewing the Markov Decision Process (MDP) [Puterman1994], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [christianos2023pangu, yang2025agentic, wang2025tutorial, wan2024alphazero], can be naturally described within this framework. Formally, an MDP is defined by the tuple
@@
where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [wang2025tutorial, wang2024openr], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [christianos2023pangu]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [wang2025tutorial]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state
@@
For large language model (LLM) agents [christianos2023pangu, wan2024alphazero, zeng2024token], the policy is typically given by the LLM itself, which maps input states or prompts directly to action distributions through its generative reasoning process:
@@ 3.2.1 Definition
At time $t$, the agent maintains episodic memory $M_{t}=\{m_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memeory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
@@
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 1) that unifies classical reinforcement learning [SuttonBarto2018] with LLM-based decision-making [christianos2023pangu], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [blundell2016model, pritzel2017neural, ramani2019short], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.

### v3

@@ Abstract
We present a theoretical study of continual and experiential learning in Large Language Model (LLM) agents that integrate episodic memory with reinforcement learning. We identify reflection, the ability of an LLM to leverage past experience to inform future actions, as the key mechanism enabling continual adaptation without fine-tuning model parameters. Empirical results on Memento [64] and Case-Based Reasoning LLMs [24, 23] show that episodic, experience-driven reflection supports generalised adaptation in a wide range of open-ended, long-horizon tasks, implying that efficient learning can occur during deployment and blurring the traditional separation between training and testing in machine learning. Motivated by this observation, we introduce the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This frames reflective memory as a control-theoretic object amenable to classical reinforcement learning analysis. We instantiate this Read–Write Reflective Learning by integrating retrieval into soft policy iteration, with guaranteed convergence. We formally show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning and retrieval-augmented generation with principled reinforcement learning and provides a rigorous mathematical basis for building reflective, memory-embedded LLM agents capable of continual general-purpose learning.
Figure 1: Learning by Stateful Reflection with Memory realises continual learning without fine-tuning LLMs by iteratively reading from (policy improvement) and writing to (policy evaluation) an evolving episodic memory that drives the agent’s online adaptation. Empirically, this Read-Write Learning paradigm has shown strong effectiveness across diverse domains, including software testing [24], automated data science [23], and deep research agents [64], demonstrating that memory-driven reflection can endow LLM agents with genuine self-improving capabilities.
@@ 1 Introduction
Modern machine learning is about learning from experience [51, 47]; both supervised learning and reinforcement learning can be viewed as processes of optimising a parameterised model through experience, either from human-labelled examples or data-driven feedback. The objective is to learn a mapping from inputs to outputs, or in the case of reinforcement learning, from states to actions, such that the resulting function generalises effectively to new, unseen situations [46, 52]. Formally, the learner assumes that the target function is constrained by a family of parameterised models, for instance, neural networks, where the parameters determine the precise form of the mapping.
@@
In supervised learning, these parameters are typically adjusted by minimising a loss function defined over labelled examples, whereas in reinforcement learning (RL), the supervision signal is indirect: instead of labelled outputs, the agent interacts with an environment, receives rewards, and adjusts its parameters to maximise the expected cumulative return [49]. In both cases, the parameters are updated by backpropagation, a gradient-based optimisation algorithm that propagates errors through the network to refine the model [43].
@@
Despite their substantial empirical successes, current machine learning paradigms suffer from fundamental inefficiencies in sample complexity. Learning via backpropagation and gradient-based optimisation typically requires an enormous number of training examples, whether labelled data in supervised learning or interaction trajectories in reinforcement learning, before competent behaviour emerges [20, 22]. From a theoretical perspective, this reflects the difficulty of identifying generalisable structure in high-dimensional, non-linear hypothesis classes when learning is driven almost entirely by statistical correlations [46]. In the absence of explicit access to semantic abstractions, compositional structure, or task-level meaning, learning algorithms must recover regularities indirectly through averaging over large numbers of samples, leading to intrinsically high sample complexity.
@@
The training of large language models continuously exemplifies this frustration. Although massive pre-training endows them with broad linguistic and reasoning abilities, adapting them to specific tasks or domains still has to resort to additional fine-tuning, such as supervised fine-tuning (SFT) or reinforcement learning with human feedback (RLHF) [37], each demanding further large quantities of curated data or human annotations [32].
@@
In contrast, humans leverage semantic representations, episodic memory, and reflective reuse of experience to adapt with far fewer examples [19]. This contrast suggests that learning mechanisms beyond purely statistical optimisation, based on stateful memory and semantic reflection, may offer a principled path to lower complexity learning without large-scale gradient updates.
@@
    In-context learning. As early as GPT-3, researchers observed that LLMs could learn from a few human-provided examples directly within a prompt, a process known as in-context learning, where a few labelled examples guide behaviour on new tasks [14] (Figure 2 (a)). For instance, given a few examples of English-to-French translations in the prompt, GPT-3 can correctly translate a new English sentence into French without any parameter updates.
@@
    Feedback-driven reflection. Subsequent studies showed that given a specific unknown task, LLMs could refine their outputs through interaction with the environment, adjusting responses based on external feedback [60] (Figure 2 (b)). For instance, in code generation, the LLM proposes a function, runs tests to observe errors or runtime feedback, and iteratively refines the code to reach a correct solution.
@@
    Internal reasoning. More recently, LLM models have displayed the ability to perform internal reasoning via chain-of-thought processes before generating final outputs [57] (Figure 2 (c)). For instance, when solving a math word problem, the model explicitly writes intermediate reasoning steps (e.g., decomposing a problem into smaller parts) before computing the final numerical answer.
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection such as information gathering or internal computation to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [14, 22], but also in multimodal LLMs [17] and non-language transformers [28, 61].
@@
In our earlier empirical work [23, 24, 64], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [23], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [64] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [24], automated data science [23, 21], and deep research [64], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
In this paper, we further provide a theoretical analysis and formalisation and leverage the capability of generalised reflection to develop a new paradigm for sample-efficient learning from experience without relying on data-intensive model training. Instead of updating model parameters, the proposed learning process operates through intermediate reflection, driven by relevant past experiences retrieved from an episodic memory [19], as shown in Fig 1. This framework can unify diverse sources of experience, whether originating from internal reasoning (e.g., chain-of-thought), interactions with the environment (e.g., ReAct [60]), or human-provided labelled data, under a single, memory-based learning mechanism.
@@
Within this formulation, learning naturally emerges from the process of retrieval: policy improvement is achieved through more effective retrieval, while policy evaluation is realised through writing experiences back to memory. This mechanism closely mirrors the principles of episodic control and reinforcement learning in humans [19, 18]. Consequently, such learning can be efficiently implemented using retrieval-augmented generation architectures equipped with large-scale memory systems [10, 31].
A key desired property is its self-learnability. With a frozen LLM [^1] While LLM parameters may evolve over time through offline updates given adequate data, such updates typically occur at a much slower timescale. Our approach does not rule out parameter learning, but instead addresses the intermediate regime in which adaptation must occur without immediate model updates. as its reasoning and reflection core, the agent improves through trial and error when data is limited and later reuses and combines past knowledge as experience grows. It balances exploitation (using known cases), exploration (targeted search in uncertain areas), and discovery (seeking new ones) through adaptive memory operations, achieving continual learning without parameter updates.
@@ 2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [59] and Graph of Thoughts [8], have demonstrated the potential of structured reflection to enhance LLM reasoning. From a different perspective, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [23, 24, 64]. To formalise this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which models continual learning as a sequence of reflective interactions with memory.
@@
Figure 1 illustrates the overall workflow of an SRDP. We begin by formally defining episodic memory as a structured repository of past interaction episodes, consistent with the classical notion of episodic memory in cognitive psychology [50].
@@
SRDP generalises and unifies several existing learning paradigms. Compared with case based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [23, 1], SRDP embeds retrieval within a sequential decision making framework. This explicitly models temporal dependencies, memory evolution, and long horizon effects under a Markovian assumption, rather than treating cases as isolated problem instances.
In contrast to retrieval augmented generation (RAG), which conditions generation on static external documents or databases [10], SRDP treats memory as a dynamic and stateful component. Episodic memory is continually updated through interaction with the environment and grounded in evaluative feedback, allowing retrieval to reflect the agent’s own experience rather than fixed external knowledge. From the perspective of memory based reinforcement learning, SRDP replaces parameter level updates with reflective memory updates mediated by an LLM. Policy improvement is achieved through structured recall and reflection over episodic memory rather than gradient based optimisation of model parameters [42]. This shift enables continual adaptation while keeping the underlying LLM fixed, positioning reflection as the primary learning mechanism.
This formulation is also strongly inspired by neuroscientific findings on the role of the hippocampus in episodic memory and reflective cognition [39, 33]. In humans and animals, episodic recall supports flexible reasoning and planning by replaying past experiences, an ability mirrored by the reflective cycles of SRDP. Thus, the Stateful Reflective Decision Process provides both a computational and cognitive bridge: it formalises how memory and reflection together enable continual, self-improving intelligence without explicit retraining.
Experimentally, it has been demonstrated that this approach enables learning without fine-tuning across various agentic tasks, including data-science agents [24], software-testing agents [23], and deep-research agents [64].
@@ 3 A Model for Reflective Process over Episodic Experience
Mathematically, the core idea is to extend the Markov Decision Process [41] with an intermediate reflection stage, where the agent retrieves relevant experiences from external memory before generating the final output, enabling continual and memory-based learning.
@@ 3.1 Preliminary: Markov Decision Process
To make self-contained, we begin by reviewing the Markov Decision Process (MDP) [41], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [16, 58, 56, 53], can be naturally described within this framework. Formally, an MDP is defined by the tuple
@@
where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [56, 55], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [16]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [56]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state
@@
For large language model (LLM) agents [16, 53, 62], the policy is typically given by the LLM itself, which maps input states or prompts directly to action distributions through its generative reasoning process:
@@ 3.2.1 Definition
At time $t$, the agent maintains episodic memory $M_{t}=\{m_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
@@
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 1) that unifies classical reinforcement learning [49] with LLM-based decision-making [16], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [9, 40, 42], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.